Archives

More... Share to Twitter Share to Facebook
News Archives

HMS Stands Up to Irene

8/28/2011

With the exception of some water intrusion and this afternoon's voltage drops experienced at the New Research Building, the Longwood Campus successfully defended itself from the driving rain and winds of Hurricane Irene. Any areas of damage or water appearing should be reported immediately to the Call Center at 617.432.1901.

 

Hurricane Irene

8/25/2011

Preparations are underway for high winds and potential for signifcant rainfall. Vanderbilt students are reminded to secure their windows. Scientists are to anticipate power interruptions. Questions should be directed to buildng managers and/or the call center at 2-1901.

 

Cooling TODAY - Friday July 22

7/22/2011

Good Morning, We are at a place at 6 am today that we may have been at twice/year in the preceding years in mid/late afternoon. We are challenged by supply vs. demand today. Know in our business as “outside design basis,” these system capabilities are not built for today’s weather. While we have managed through this in the past, never have we been confronted with these conditions so early in the day without relief expected later in the course of the day. Please have patience, and do whatever you can to reduce heat load by closing window blinds, turning off lighting, deferring operation of heat generating equipment, turning off equipment not in use, limiting travel in/out of cold rooms and keeping doors and windows closed. Saturday will find us better off with continuing improvement throughout the weekend and next week.

 

Heat & Humidity

7/19/2011

Once again, the New England version of tropical weather descends on us. Your cooperation in conserving cooling and electricity usage this week will assist in ensuring uninterrupted supply. Keeping windows closed and shades drawn can make an impact. Questions/suggestions can be directed to facilities@hms.harvard.edu

 

Facility Contractor Change

6/29/2011

With high expectations and enthusiasm, HMS welcomes UGL Services TODAY as the Facility Maintenance & Operations Contractor. Please pay particular attention to personnel assignment changes and new contact information. Service requests now via 2-1901 OR facilitiescallcenter@hms.harvard.edu

 

Heat Extreme & Subsequent Storm

6/10/2011

High heat and humidity arrived in the past 72 hours. Systems performed admirably, chilled water supply temperatures approached their maximum value as did supply tonnage. Relief in the early evening Thursday afforded recovery. Thanks to all for the conservation initiatives.

 

Heating to Cooling Swap

4/26/2011

Those buildings services with a single feed line will be converted from hot water to chilled water beginning Wednesday, April 27th. Space will be controlled by cooling only until the Fall.

 

Earth Week has Arrived

4/18/2011

Join in on our post consumer composting launch this week in the Atrium and Courtyard cafe, attend the Green Lab Lecture, Tuesday, Armenise Amphitheater, 5:30-7:30pm, the HSPH Faculty Panel, Thursday, Kresge G-3, 12:30-1:45pm and our bike maintenance clinic April 25th, Armenise Amphitheater, 12:30-1:30pm.

 

aaahhhh ! A Taste of Spring

3/18/2011

After a snowy winter. Beautiful 2 days in Boston - temperatures in the 60F's with blue sky and sunshine !!!!!

 

Natural Gas Supply - LHRRB & Mudd

3/4/2011

This weekend, a wicking system will be installed by the Gas Company to "spike" our gas supply with mercaptan at the point of delivery to these buildings. An undetectable amount of added odor to the gas has caused us to interrupt that service until this system can be installed. Clearly, undetectable gas can lead to a dangerous situation from leaks or inadvertent release.

 

Building & Garage Access TODAY

2/17/2011

We are experiencing a campus wide problem with access this am. Please be patient as personnel may not get customary access to garages or all entrances. This is random, not everyone is affected. Security is working to resolve this issue.

 

JCI Celebrates 2 Year Safety Record

2/9/2011

Earlier today, JCI celebrated their 2 year record of consecutive days without an OSHA recordable or lost time incident,in a luncheon with their HMS Campus employees. The JCI record continues at 740 days today. CONGRATULATIONS JCI !!!!!!!!!

 

Goldenson I Bike Cage Roof Collapse

2/2/2011

As a direct result of the excessive snow followed by rain, we have had a structural failure of the roof. Bike owners are asked to contact Security for instruction regarding their property. Access is prohibited until the roof has been safely removed, expected some time later this evening.

 

Temperature Extreme - Steam Usage

1/20/2011

Temperatures expected this Saturday through Monday will call for record usage of steam Monday morning unless we take some mitigating actions. Between the hours of 6 am and 9:30 am on Monday January 24th, we ask that staggered schedules for cage/glass washing, and sterilizers be implemented so as to avoid a spike (and $$ demand charge) in steam required. Any questions please call Allen Hebert 617.432.7997

 

Significant Snowstorm

1/12/2011

Another substantial snowfall at Longwood Ave this season. Snow has high moisture content in depths expected to be 15+" calling for our raising temperatures 10+F in Courtyard Cafe,TMEC Atrium and Gordon Hall Atrium to warm the skylight glass facilitating snow melt and snow load reduction. So far so good. Snow removal crews headed by Bob Christiano doing a remarkable job keeping passageways open.

 

2011 A New Year - Some Must Wins

1/7/2011

As we envision our expected accomplishments for this coming year, our focus is drawn to three key areas 1) LCME - Med School Accreditation 2) Implementation of new Energy Conservation Measures in our march to reducing GHG emissions and 3) supporting a massive change related to science department space planning within the facility.

 

HSDM Main -Small Fire

12/10/2010

A small electrical fire in the lower level has disabled the vacuum system. While little to no damage to the facility occurred, patient appointments this morning have been cancelled with effort underway to restore the system for afternoon patient service. An update will be provided later this morning.

 

HSDM Main - UPDATE 11:15 am

12/10/2010

A replacement vacuum motor has arrived. It is expected that the system will be returned to service by 1 pm today.

 

A Word About Space Heaters

11/5/2010

Space heaters are NOT permitted at HMS unless approved and registered by HMS Facilities. Requests for approval should be made through the respective Building Manager.

 

HVAC Isolation Test - Started

10/21/2010

 

HVAC Isolation Test Concluded

10/21/2010

 

All work safe to resume

HVAC Isolation Test Oct 21 Noon to 1 pm

10/20/2010

The Facilities Group will be testing an emergency shutdown feature. Shutdown requires that fume hood work cease and the sash be closed. Other precautions include refraining from using doorways in to/out of Vivariums, and the management of steam escape from autoclave and glasswasher doors during this period. Questions ? Now - Robert_beherent @hms.harvard.edu, 2-1429 During the Test - 617.432.4873

 

Fall Semi-Annual Freecycle - Wed Oct 20, 9-1

9/22/2010

DATE CHANGED ! Allow others to reuse what you don't need ! Donations will be accepted leading up to the event in the TMEC Building, Room 150. The room is open for drop-off specific hours to follow Office supplies and housewares are welcome (see a full list of acceptable donations here); for furniture, large electronics and other larger items please use the Harvard ReuseList http://green.harvard.edu/reuselist, a collaborative effort between the Office for Sustainability and Harvard’s Office for Strategic Procurement to facilitate the freecycling of resources within the Harvard community. Questions? Please email Rachel Johnson at Harvard’s Office for Sustainability at rachel.johnson@post.harvard.edu.

 

Earl is a Dud...Dissimilar Metals Get Us

9/8/2010

While Hurricane Earl passed us by with minimal rainfall, we experienced a chilled water piping failure in Gordon Hall specifically related to a brass to carbon steel threaded joint. Damage on the 3rd and 2nd floor is being remediated. A plan to complete an entire building inspection is in development.

 

Hurricane Earl

9/3/2010

Preparations for the arrival of Earl have been underway since Tuesday. Precautions have been taken, contingencies in place, and staffing adjusted to better coincide with the expected arrival tonight and very early morning tomorrow. Forecasters now paint a much improved picture from mid-week for the immediate Boston area. As always, report any issues to the onsite Call Center at 617.432.1901.

 

HMS Facilities Personnel Change

9/2/2010

We are pleased to welcome Sharon Lorince to HMS Facilities as an Associate Director of Facilities. Sharon is joining us from Harvard FAS on Tuesday September 07. Sharon will be primarily responsible for activities north of Longwood Avenue, but not exclusively. We are excited to have a new team in place next week and look for the advance of continuous improvement.

 

HMS Facilities Personnel Changes

8/20/2010

We are fortunate to announce that John Paras and Rob Behrent will be joining the HMS Facilities Organization in the coming weeks. John as an Associate Director of Facilities and Rob as the Facilities Engineer. Each of these individuals brings a wealth of talent and experience to our organization. Please join me in welcoming them.

 

Another Round - Intense Rain Storm

8/5/2010

Once again, late this afternoon, we have experienced another deluge of rain affecting the Longwood Campus. We have first hand visual knowledge now, of backup from the city system into ours. We will be addressing this with MASCO on Aug 06. Areas affected are Gordon Hall basement and Mudd vivarium. Water and damage in no way resembles the magnitude of the last occurrence. Our efforts over the past weeks in cleaning and underground camera inspection provides us comfort in knowing our HMS systems are performing as intended.

 

Flash Flooding at HMS

7/11/2010

Yesterday we experienced some affects of the short duration but intense rainfall in late afternoon. Areas most affected are the HIM & Mudd vivariums. Animal care management was on campus late last night. Other areas experienced some water infiltration. In all but one case these areas were impacted by a backup in drainage experienced as well throughout the City. Affected groups have been notified, some via early morning email today. Drying and dehumidification has been ongoing since late night/early morning. Concerns should be brought to the Building Manager.

 

Extremes - Heat & Humidity - Over for This Week

7/9/2010

The week long heat and humidity extremes have come to a close. To those who were willing and able to reduce building heat loads. Thank you, you had an impact. We managed our way through the first two levels of a four level alert system from the local power plant and avoided extraordinary curtailment on the Longwood Campus. Normal operations can safely resume. In areas where you found that conservation measures had insignificant impact on your activity, let’s look to integrate that into our daily routine by contacting your respective Building Manager. Thank you again.

 

Building Cooling & Heat Loads

7/6/2010

Today and the next few days this week will present challenges to the HMS Building Operating staff. Please make efforts to conserve heat generation including reduced lighting needs. Where some are fortunate enough to have taken vacation this week, let’s ensure their areas are as energy efficient as possible in their absence. All issues should be reported to the Call Center at 2-1901.

 

Cooling Restrictions

6/28/2010

The extreme heat, humidity and elevated chilled water supply temperatures are causing difficulty in maintaining targeted cooling levels. Conservation including reducing heat generating sources is requested. Relief is expected this afternoon with the arrival of a cool front through the Boston area.

 

HMS Facility Engineer

6/18/2010

It is with disappointment that I announce that our valued engineer, Brian Barmmer, has elected to leave HMS and move on with his career elsewhere. Join me in congratulating Brian in his new position and wish him well. Recruiting for replacement is imminent.

 

Water Dispensing Station - TMEC

5/28/2010

The system became operational this week. Hot & Cold, filtered city water is now available for consumption on the 2nd floor in the vending area. LED lighting illuminates the station. Thanks go to Students for Environmental Awareness in Medicine (SEAM) for their perseverance in moving our culture away from bottles and bottled water and taking advantage of the fluoride that city water affords.

 

Commencement 2010

5/27/2010

Congratulations to students and families on this very important day and continued success in their ventures forward. Our best to all !

 

WAB Fire & Water Damage May 07

5/14/2010

An unfortunate early morning fire on the 3rd floor was extinguished with minor facility fire damage. The ensuing sprinkler discharge estimated at 2200 - 2500 gallons travelled to the basement level affecting the south end of the building. Standing water clean up was rapidly completed, measurements were taken for wall and ceiling moisture, walls/ceilings removed where necessary and a thorough drying program initiated that same day into May 08,09,10. As of Tuesday, May 11 restoration of the facility began. The apparent cause resides with electronics in an idle experiment.

 

MWRA Water Supply - Resolved

5/4/2010

Early Tuesday morning, MWRA and the Governor's Office removed the boil water order for Boston. MWRA recommends that all water lines be flushed prior to return to normal use. HMS and JCI are in the process of flushing all fixtures and expect to have the campus returned to normal operation by early Wednesday morning.

 

MWRA Water Supply - Boil Order

5/2/2010

Following notification mid afternoon Sat May 01, dormitory residents were notified and all Longwood Campus HMS buildings posted. A plan was set in motion for animal care. Ice machines were isolated from the water supply. Many activities expected to be impacted Monday May 03.

 

Greenhouse Gas Reduction - Moving Forward

4/26/2010

This week marks the kickoff of the 1st formal recommissioning project or Energy Conservation Method (ECM),in the Warren Alpert Building. Recommissioning is expected to be the most influential activity in reducing greenhouse gas - energy usage equivalent on the Longwood Campus. Given the laboratory complexity and building size, this is expected to continue for the next few months. Allen Hebert, HMS Energy Specialist is spearheading this activity.

 

Concluding Earth Week

4/23/2010

What a week! Thanks to the Office for Sustainability, Students for Environmental Awareness in Medicine (SEAM), HMS Operations, interested and enthusiastic vendors. With fairs held in the NRB Tuesday, "FLOW" shown Wednesday and the 2nd fair in the Courtyard Cafe, along with free bicycle tune ups Thursday, the celebaration concludes TODAY with "Green Carpet Awards" in Sanders Theater later this afternoon. Congratulations to all HMS award recipients and those nominated for their efforts in seeing to a more sustainable Medical School campus.

 

Earth Week 2010

4/19/2010

Join us for Earth Week activities on the Harvard Longwood Campus this week. Please follow this link for the "HMS Only" schedule.... HMS_Earthweek_Events.pdf

 

Record Heat

4/8/2010

High temperature record set in Boston yesterday at 90F. Medical School Longwood campus at 91F. Single feed systems remain on hot water/steam with target transition to chilled water on April 15th. Temperatures expected to moderate today to more seasonable levels.

 

Spring Arrives

4/2/2010

Following days of torential rain and widespread area flooding,(some reporting as a 200 year storm)rivers have reached crest and are returning to their embankments, damage assessment and recovery begins, for some, the second time within a month. HMS Longwood battles dozens of water intrusion occurences while basking in blue sky, sunshine and temperatures expected to reach 70s today and the upcoming weekend.

 

HSDM Main - Elevator

3/24/2010

Work continues to replace the jacking system for the elevator that delivers patients to the clinic. This elevator was placed Out of Service the middle of this month from a hydraulic leak, suspect of growing. Today, remediation of the area beneath the jack should complete. The new components are due to arrive within the next 10 days. Patients are directed up the Longwood staircase or to the REB for same level access to the clinical area.

 

Record Setting Rainfall

3/16/2010

Enduring three consecutive days of heavy rains has caused numerous water intrusion issues throughout the campus from water travelling through exterior doorways to wind driven rain around windows. All locations of water intrusion must be reported to the Facilities Group at 2-1901 for response, tracking and evaluation. Granted, this has been unusual, repairs and design improvements are a must going forward to preclude as much as possible from again affecting HMS in the future.

 

C Building Odor

3/16/2010

Very early this morning an elevator component overheated. While work continues to determine the cause and make suitable repairs, a residual odor lingers within the building. This odor should not be a concern for the occupants. Any questions, please contact the Call Center at 2-1901.

 

Energy Display Recognized

3/12/2010

The HMS Interactive Energy Display was recognized this week in the HMS publication, "Focus", March 5 edition and the International Facility Management Associations (IFMA) "Facility Management Journal",Industry News, March/April edition.

 

WAB Garage Odor

2/17/2010

The foul odor detected in the WAB parking garage was attributed to a grossly failed battery trickle charge system. The batteries are an integral part of the "stacker' system. The odor was properly dealt with by 10 am yesterday, February 16th.

 

HMS Temperature Policy

2/12/2010

The HMS Executive Dean for Adminstration has formally released the subject document to the HMS Longwood Community. This Policy dovetails into the University Policy published last Summer. The Facilities Group objective is to deliver HVAC service to all areas as thoughtfully & consistently as possible in accordance with this Policy. Our collective efforts to conserve energy usage has a direct bearing on reducing greenhouse gas emmissions. As always, we appreciate hearing of performance issues from you. While we take great care to monitor these systems and components your input provides us with more data, instrumental in troubleshooting an issue.

 

SEAM Leadership Focuses on Bottled Water

1/29/2010

Students for Environmental Awareness in Medicine have taken serious interest in changing Student, Staff and Faculty consumption of plastic water bottles and delivered , flouride free, purified water. First up, a look at TMEC in concert with the HMS Facilities Group to seek out potential locations for personal bottle/cup fillers using Boston City Water. Currently under consideration, the 2nd floor atrium, vending machine area... more to come

 

MATEP Chilled Water to HIM - Restored

1/12/2010

Following the isolation of a seasonal branch line leak to 333 Longwood Ave., MATEP restored chilled water service to the HIM on Friday, January 8, 2010 @ 8:55 am.

 

MATEP Steam & Chilled Water Disruption

1/8/2010

At approx 2am Thursday, Jan 07 HMS experienced a significant loss in steam pressure throughout the campus. MATEP removed the "B" header from service believing a significant steam leak was underway at Binney St & Longwood Ave. Steam pressures were restored about 5am when it was determined that the leak was chilled water but vaporizing due to the proximity of their steam lines to the leaking chilled water. The HIM was isolated from the system, the balance of campus continues operating normally. The HIM remains isolated without chilled water as of 7am Friday, Jan 08. MATEP has discovered a split line into the Harvard Coop @ 333 Longwood. MATEP believes that line can be isloated and that BIDMC, Children's and HIM will have chilled water restored during the day today. The HIM temperatures remain stable. In much of the HIM, 100% outdoor is required and conditioned before distribution into the building. Outdoor temperatures remain cold enough that only two equipment rooms require supplemental cooling and have been functioning adequately over the past 24 hour period.

 

Gas Delivery

11/19/2009

It was recently discovered that our supply of gas to HMS Longwood was odorless at the lab level. We were educated about "odor fade" by National Grid. Since that time the addition of a Mercaptan Blend has been being administered locally by National Grid to purge our lines with this odor additive. The Mercaptan add is nearly complete and gas odor detection restored.

 

TMEC Electrical Fire Early AM Thursday Sept 17

9/17/2009

We experienced a small electrical fire in the TMEC originating from the Fire Pump, Motor Control Panel. Boston Fire responded. A contract worker was transported to Mass Eye & Ear, not sustaining any apparent serious injury. There is some residual smoke smell within the building, not of concern for occupants. Power was restored at approimately 5 am.

 

Summer 2009 - It's Over

9/3/2009

A good healthy 3 weeks of warm weather and some humidity has left us all awakening to cool morning temperatures and wishing for the summer that wasn't.

 

Weather - Summer 2009

8/6/2009

The heat and humidity arrived at the last minute in July. Savings in energy for June/July is appreciable

 

Weather

6/30/2009

Miserable, cool, rainy June 2009. Saving on chilled water usage - extended shoulder season

 

Commencement 2009

6/4/2009

A stunning day ! Our thoughts with all who move on. Best wishes.

 

2009 Medical School Commencement

6/2/2009

June 4th, tent is up

 

Welcome To Our New Website

5/13/2009

Our website had been redesign designed to make it easier for customers to access needed facilities information